OpenStudy (anonymous):

why does [ limit (x to 10) ] give x = infinity ??

OpenStudy (anonymous):

does it ?

OpenStudy (anonymous):

\[\lim_{x\rightarrow 10}x = 10\]

OpenStudy (anonymous):

it does, source : http://www.wolframalpha.com/input/?i=limit+%28x+to+10%29 unless this is wrong.

OpenStudy (amistre64):

\[\lim_{x->10^+}\frac{1}{(x-10)}=\infty\]

OpenStudy (amistre64):

lol ... um thats just the wolf giving you the best interpretation it knows how with regards to what you inputed

OpenStudy (amistre64):

in other words, the wolf doesnt like to say, " i dont know" and starts making up answers to solve imaginary questions lol

OpenStudy (anonymous):

ahhhh! so it can be utter uslessness ! I See

myininaya (myininaya):

\[\lim_{x \rightarrow 10}x=10\]